OpenStudy (anonymous):

Does anyone know how to find the maximum value for y=-2x^2+8x-5?

jimthompson5910 (jim_thompson5910):

You need to find the vertex. Do you know how to do this?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

What do you get for the vertex?

OpenStudy (anonymous):

-b/2a

jimthompson5910 (jim_thompson5910):

in this case, a = -2 b = 8

OpenStudy (anonymous):

I mean 2

OpenStudy (anonymous):

sorry, my bad

jimthompson5910 (jim_thompson5910):

So what is the x coordinate of the vertex?

OpenStudy (anonymous):

I"m not sure

jimthompson5910 (jim_thompson5910):

Plug in a = -2 and b = 8

OpenStudy (anonymous):

wait into -b/2a?

jimthompson5910 (jim_thompson5910):

Yes, into -b/(2a)

jimthompson5910 (jim_thompson5910):

x = -b/(2a)

OpenStudy (anonymous):

oh, I did I received 2 ( I mentioned in the earlier comments)

OpenStudy (anonymous):

x=2

jimthompson5910 (jim_thompson5910):

good, that's the x coordinate of the vertex

jimthompson5910 (jim_thompson5910):

Plug that x value into y=-2x^2+8x-5 to get the y coordinate of the vertex

OpenStudy (anonymous):

wait is that all?

OpenStudy (anonymous):

but isn't that how u find the minimum?

jimthompson5910 (jim_thompson5910):

That's if a > 0 In this case, a < 0

jimthompson5910 (jim_thompson5910):

If a > 0, then y = ax^2 + bx + c will have a minimum If a < 0, then y = ax^2 + bx + c will have a maximum

OpenStudy (anonymous):

ohh, so technically finding the minimum and maximum can be solved by the same way?

OpenStudy (anonymous):

well, ok thank-you I understand it now

jimthompson5910 (jim_thompson5910):

Yes they are solved the exact same way. The only difference will depend on the value of 'a' as shown above.

OpenStudy (anonymous):

ohhhhhhhhhhhhhh, ok thank-you again

jimthompson5910 (jim_thompson5910):

yw